OpenStudy (anonymous):

6 over the cubed square root of 16

OpenStudy (anonymous):

\[\frac{ 6 }{ \sqrt[3]{16}}\]

OpenStudy (anonymous):

If I understand your question you need to rationalize the denominator! So we multiply both numerator and denominator by Cube root of 16. We can do this because cube root of 16 over cube root of 16 is 1, and multiplying anything by 1 doesn't change the value so... \[\frac{ 6 }{ \sqrt[3]{16} } * \frac{ \sqrt[3]{16} }{ \sqrt[3]{16} } * \frac{ \sqrt[3]{16} }{ \sqrt[3]{16} }\] ------ We end up with: \[\frac{ 6 * \sqrt[3]{256} }{ 16 }\] ---- Simplify \[\frac{ 3\sqrt[3]{4} }{ 2}\]

OpenStudy (anonymous):

Let me double check this though... been a long time.

OpenStudy (anonymous):

That's it, whew. Any questions please let me know!

OpenStudy (anonymous):

thats not right either

OpenStudy (bobo-i-bo):

Do you know your laws of radicals?

OpenStudy (anonymous):

please refresh my memory

jimthompson5910 (jim_thompson5910):

@Dangazzm has the right answer. It could be that either the computer system is glitching or you might have typed in the answer incorrectly If you can, please post a screenshot of what you typed in

OpenStudy (anonymous):

this is what i typed

OpenStudy (anonymous):